OpenStudy (anonymous):

49 to the negative 1 half power.

OpenStudy (jdoe0001):

noted, thanks for sharing

OpenStudy (anonymous):

huh????

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

hint: x^(-1/2) = 1/x^(1/2) = 1/sqrt(x)

OpenStudy (anonymous):

is the answer 7???

Vocaloid (vocaloid):

not quite x = 49 1/sqrt(49) = ?

OpenStudy (anonymous):

1 over 7 ?

Vocaloid (vocaloid):

yea

OpenStudy (anonymous):

okay thank you